John is 6 years older than his brother. He will be twice as old as his brother in 4 years. How old is John now?
Answer by mananth(16949)

brother's age x
John's age x+6
4 years later
brother's age x+4
John's age x+6+4=x +10
Jack is 2 times older
x+10=2(x+4)
x+10=2x +8
2x-1x=10-8
x=2
/1
x=2 brother's age
x+6 =8years John's age